# Detection of ChatGPT Fake Science with the xFakeSci Learning Algorithm

### v3

@@ 4.3.1 Model Derivation and Network Construction
Since our training model is a network driven from text, we used the Term Frequency-Inverse Document Frequency (TF-IDF) [40, 41, 42, 43, 44, 45] to extract word features as building blocks of the training models. The TF-IDF algorithm can be configured to generate two consecutive words (known as bigrams) that may prove significant across an entire dataset [46, 47]. Equation 2 shows the mathematical representation of the TF-IDF. where $\text{tf}(t,d)$ is the frequency of bigram $t$ in document $d$,$\text{idf}(t,D)$ is the inverse document frequency of bigram $t$ in the document set $D$.
@@
Algorithm 2 outlines the steps involved in building the network model from bigrams. We applied this algorithm twice, once to create a publication training model and another to create a ChatGPT model. We recorded the corresponding statistics (number of nodes and edges) for each model in Table 2. The initial observations revealed a consistent pattern where models constructed from ChatGPT-generated text exhibited the lowest number of nodes, yet they also maintained the highest number of edges. The resulting models showed disconnected components and fragmented communities, requiring pruning. This need was satisfied by applying the Largest Connected Components (LCC) algorithm [48], which ensured that the resulting networks maintained high connectivity. The LCC presents an admissible pruning heuristic due to the presence of high-degree nodes that promote network stability and robustness [49, 50, 51, 52].
@@ 4.4 Evaluating the Premise of ChatGPT’s Distinctive Behavior
We summarized the analysis of each dataset and disease in Table 3. The initial observations indicated that the ChatGPT ratios fluctuated between 27% and 30% for Alzheimer’s disease, 27% and 29% for cancer, and 28% and 32% for depression. In contrast, the ratios derived from scientific articles ranged between 14% and 16% for Alzheimer’s disease, 14% and 17% for cancer, and 9% and 11% for depression. The full analysis of these results will be discussed in the “Results” section. The analysis demonstrated that the ratios of ChatGPT-generated documents were significantly different from those computed from scientific articles. This distinction serves as proof that the premise is indeed true. Furthermore, this knowledge provides lower and upper bounds for each disease, offering more guidance to the algorithm to predict the label while avoiding issues of overfitting. Algorithm 4 demonstrates how the ratios are computed. For brevity, we only demonstrate for the depression disease. Table 4 presents the corresponding lower and upper bounds, which are necessary during the calibration phase.
@@ 4.4.1 Label Prediction of Articles: Real vs Fake
Testing the premise in the section above demonstrated the fundamental differences in content behavior between fake ChatGPT articles and real publications. In this section, we present the learning algorithm, which is the main contribution of this paper. During the Coronavirus global pandemic, our previous work addressed the challenge of detecting fake news and science as an emerging infodemic [15]. However, this work was limited by the lack of comprehensive machine-generated datasets that could adequately assess performance in the presence of fake data. Now, with the advent of ChatGPT and generative AI technologies, we can generate diverse datasets using prompt-engineering algorithms as demonstrated above. Additionally, the previous work did not make use of data-driven insights, which we incorporate as a calibration step. This is an intermediate phase that takes place after the training phase and before the label prediction phase. Due to these factors, the previous work was limited to a single-mode label prediction using a single type of dataset. Therefore, it was necessary to split the dataset into training and test sets. The following is a complete comparison of the previous work and the current features presented by xFakeSci, such as content type, configuration parameters, classification mode, calibration, and classification, presented in Table 1.
@@
Content type publication/news publications/ChatGPT fake article
@@
Classification/Year not-supported compares classification by the year of publication
@@
As shown in Table 1, the xFakeSci algorithm is particularly designed to address multi-mode classification. Therefore, it is expected to train the algorithm using two or more independent types of data. Consequently, the algorithm also expects a hybrid test set of mixed types and will produce more accurate labels for each type. However, such modes suffer from what is known as the “overfitting” issue [53, 54]. The introduction of the calibration step (by calculating the lower/upper bound ratios captured in Table 4) was to guide the decision of the final label prediction and avoid such an issue. The table demonstrates the clear separation of lower/upper bound ratios. Therefore, we further utilize such a mechanism by incorporating a calibrating step to further guide the classification process without having to train the algorithm with too many samples. The algorithmic steps for the calibration process are explained in Algorithm 4. Though the ranges provide an extra net for predicting the label, it is also possible that some document instances may fall outside the specified ranges of the datasets, which could result in not predicting a label. Therefore, we introduced a proximity heuristic that favors the shortest distance to the ranges driven from the individual datasets (real or ChatGPT-generated) and assigns a label accordingly. Equations 5 and 6 demonstrate how the distance is calculated.
@@
Algorithm 5 illustrates the computational steps for multi-mode execution, demonstrating the complexity involved, including the proximity distance. To use the algorithm in detecting fake science, it must be trained using two different types of data: (1) a real publications dataset and (2) ChatGPT-generated datasets. The algorithm also expects the ratio means of each data source, which are computed using the calibration algorithm.
@@
3:Input: a dataset $D$ of mixed documented fake and real
@@ 5 Discussion
In a world where Generative AI has become widespread, our study aimed to investigate the potential issues of using ChatGPT to generate fake science. Specifically, we first tested the intuition of whether the content generated by ChatGPT may exhibit unique characteristics that distinguish it from real research. We explored this task using prompt-engineering, where we created engineered datasets on the subjects of Alzheimer’s, cancer, and depression diseases. In this work, we contributed a prompt-engineering algorithm on how to generate simulated content to evaluate this premise. Working with plain text (using publication Abstracts or generated from ChatGPT), using the TF-IDF algorithm is a common approach to generate bigrams that can be used to construct more complex models.
@@
Our initial observation of networks generated from ChatGPT content is that they are highly connected and contain fewer nodes compared to networks constructed from publication text. Additionally, when we calculated the ratios of the number of bigrams against the total number of words of documents on k-Folds, we found that the ratios of ChatGPT content are much higher than scientific abstracts. These two indications supported our intuition that ChatGPT documents exhibit distinguishable behavior than PubMed abstracts. One interpretation of this observation could be due to the inherent design of the ChatGPT engine. As observed, ChatGPT is optimized to generate highly convincing content by predicting the next correlated terms statistically using a Large Language Model. On the other hand, scientists prioritize accurate documentation of hypotheses testing, scientific experiments, and careful explanation of observations. Describing science in terms of highly correlated words is not a goal of scientists. Clearly, the difference in goals may contribute to less connectivity in scientific publications.
@@
Further, we introduced the xFakeSci algorithm, a learning algorithm that predicts a label for a given article. In the Methods section, we showed that it is designed to operate in two modes: (1) Single-mode: where only one type of article from the same source is used for training and a new set of documents from the same pool is used for predicting the label of an article; (2) Multi-mode: where the algorithm was trained from two sources and a hybrid train model (of real and generated datasets) was constructed to make the predictions. The single-model is trivial; therefore, we focused our experiments on demonstrating the multi-mode. We performed several experiments to do the following: (1) to test and measure how the xFakeSci algorithm predicts labels of ChatGPT generated documents for a given disease when mixed with scientific abstracts, (2) evaluated whether the algorithm performs consistently using various datasets of different diseases not only one disease, (3) to test whether the year of the publication plays a role in predicting ChatGPT generated documents when mixed with publications from various periods (2020-2024, 2015-2019, 2010-2014), (4) benchmark the algorithm against the state-of-the-art (SOTA) algorithms. Our results for each experiment used the TP, TN, FP, FN metrics and F1 scores.
@@
When testing whether the year of publication plays a role in label prediction, we observed F1 scores of 91%, 92%, and 94% for cancer-related publications across different periods. This suggests a pattern of better detection of ChatGPT articles when mixed with older publications. However, identifying newer publications proved more challenging. For Alzheimer’s disease, while no improvement was observed, degradation was also absent. As mentioned earlier, the Alzheimer’s datasets were the smallest among all other datasets, limiting the calibration process due to fewer k-Folds compared to other diseases. In the case of depression, the algorithm exhibited consistent performance with an F1 score of 80% across all periods. It’s plausible that mental health data acquisition posed limitations, potentially constraining resources from this specific area. Testing this hypothesis involves measuring document similarity between PubMed and ChatGPT sources using lexical and semantic analysis.
@@
Upon benchmarking xFakeSci against some of the state-of-the-art (SOTA) algorithms, we observed an interesting pattern: xFakeSci correctly predicted all the scientific publications in all the experiments we performed. However, other algorithms misclassified publications as ChatGPT and vice versa (true positives, false positives, false negatives, and true negatives). xFakeSci, however, needed improvement in predicting true negatives (ChatGPT documents), as many ChatGPT documents were labeled as true positives (real publications). In all the experiments, the F1 scores of xFakeSci ranged between 80% and 94%. In contrast, the SOTA algorithms showed much lower performance, with F1 scores ranging between 32% and 52%. We attribute the high performance of xFakeSci to the calibration process, which was guided by ratios and proximity distances. Although the training model remained lightweight, both heuristics provided more guidance for predicting fake articles. This novel calibration method benefits from an abundance of data, without suffering from overfitting issues like other SOTA classification algorithms.
@@ 6 Conclusions and Future Directions
When we asked a high school student about their knowledge of ChatGPT, they responded, “Do you mean that tool that does my homework for you?” Indeed, ChatGPT is an incredibly sophisticated tool with a wide range of impressive capabilities. Since the rise of ChatGPT, many new research avenues have opened, and many long-standing questions are now being investigated. However, the most significant concern associated with ChatGPT and other generative AI tools is that they could pose a threat to the future of science. If younger generations utilize ChatGPT to plagiarize, it could undermine the integrity of research and learning, potentially having a negative impact on the development of future pioneers.
@@
Looking ahead, there are several avenues for future research based on our current work: (1) conducting a preprocessing step (e.g., clustering) to group more closely related publications together (e.g., breast cancer, prostate cancer, and others), or separate diseases from co-morbidities. The use of knowledge graphs may be a powerful tool to use in continuing to investigate this research direction; (2) further experimentation in training and calibrating the xFakeSci algorithm by utilizing heuristics learned from preprocessing steps and the discoveries of clusters; (3) testing the algorithm on more than two data sources (clinical reports, publications, and ChatGPT-generated).
@@ Acknowledgments
This research is supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement Sano No 857533 and carried out within the International Research Agendas programme of the Foundation for Polish Science, co-financed by the European Union under the European Regional Development Fund, Additionally, is created as part of the Ministry of Science and Higher Education’s initiative to support the activities of Excellence Centers established in Poland under the Horizon 2020 program based on the agreement No “MEiN/2023/DIR/3796”, and and the National Natural Science Foundation of China (NSFC) under grant 62120106008. The authors also acknowledge Laila Hamed for her valuable perspective on ChatGPT.
@@
A.H. conceived the idea(s), A.A. and X.U. designed the experiment(s), A.A. and X.W. analyzed the results. All authors written and reviewed the manuscript.
@@
Table 6: Multi-Mode Experiments: F1 Classification Scores By Periods
@@
Figure 4: Multi-Mode Experiments: F1 Classification Scores for (Cancer, Alzheimer’s, and Depression) for publications gathered in period (2019-2024.)
Figure 5: Showing (TP, TN, FP, FN) Mertics and F1 Scores for xFakeSci vs SOTA algorithms Using Cancer Dataset From 2020-2024
Figure 6: Accuracy of xFakeSci vs SOTA Showing Publication Year as a Prediction Factor.

### v4

@@ 4.3.1 Model Derivation and Network Construction
Since our training model is network driven from text, we used the Term Frequency-Inverse Document Frequency (TF-IDF) [40, 41, 42, 43, 44, 45] to extract word features as building blocks of the training models. The TF-IDF algorithm can be configured to generate two consecutive words (known as bigrams) that may prove significant across an entire dataset [46, 47]. Equation 2 shows the mathematical representation of the TF-IDF. where $\text{tf}(t,d)$ is the frequency of bigram $t$ in document $d$,and $\text{idf}(t,D)$ is the inverse document frequency of bigram $t$ in the document set $D$.
@@
Algorithm 2 outlines the steps involved in building the network model from bigrams. We applied this algorithm twice, once to create a publication training model and another to create a ChatGPT model. We recorded the corresponding statistics (numbers of nodes and edges) for each model in Table 3. The initial observations revealed a consistent pattern where models constructed from ChatGPT-generated text exhibited the lowest number of nodes, yet they also maintained the highest number of edges. The resulting models showed disconnected components and fragmented communities, requiring pruning. This need was satisfied by applying the Largest Connected Components (LCC) algorithm [48], which ensured that the resulting networks maintained high connectivity. The LCC presents an admissible pruning heuristic due to the presence of high-degree nodes that promote network stability and robustness [49, 50, 51, 52].
@@ 4.4 Evaluating the Premise of ChatGPT’s Distinctive Behavior
We summarized the analysis of each dataset and disease in Table 4. The initial observations indicated that the ChatGPT ratios fluctuated between 27% and 30% for the Alzheimer’s disease, 27% and 29% for cancer, and 28% and 32% for depression. In contrast, the ratios derived from scientific articles ranged between 14% and 16% for the Alzheimer’s disease, 14% and 17% for cancer, and 9% and 11% for depression. The full analysis of these results will be discussed in the “Results” section. The analysis demonstrated that the ratios of ChatGPT-generated documents were significantly different from those computed from scientific articles. This distinction serves as proof that the premise is indeed true. Furthermore, this knowledge provides lower and upper bounds for each disease, offering more guidance to the algorithm to predict the label while avoiding issues of overfitting. Algorithm 4 demonstrates how the ratios are computed. For brevity, we only demonstrate for the depression disease. Table 5 presents the corresponding lower and upper bounds, which are necessary during the calibration phase.
@@ 4.4.1 Label Prediction of Articles: Real vs Fake
Testing the premise in the above section demonstrated the fundamental differences in content behavior between fake ChatGPT articles and real publications. In this section, we present the learning algorithm, which is the main contribution of this paper. During the Coronavirus global pandemic, our previous work addressed the challenge of detecting fake news and science as an emerging infodemic [15]. However, this work was limited by the lack of comprehensive machine-generated datasets that could adequately assess performance in the presence of fake data. Now, with the advent of ChatGPT and generative AI technologies, we can generate diverse datasets using prompt-engineering algorithms as demonstrated above. Additionally, the previous work did not make use of data-driven insights, which we incorporate as a calibration step. This is an intermediate phase that takes place after the training phase and before the label prediction phase. Due to these factors, the previous work was limited to a single-mode label prediction using a single type of dataset. Therefore, it was necessary to split the dataset into training and test sets. The following is a complete comparison of the previous work and the current features presented by xFakeSci, such as content type, configuration parameters, classification mode, calibration, and classification, as presented in Table 2.
@@
Content type publication/news publications/ChatGPT fake articles
@@
Classification/Year not-supported comparative classification by the year of publication
@@
As shown in Table 2, the xFakeSci algorithm is particularly designed to address multi-mode classification. Therefore, it is expected to train the algorithm using two or more independent types of data. Consequently, the algorithm also expects a hybrid test set of mixed types and will produce more accurate labels for each type. However, such modes suffer from what is known as the “overfitting” issue [53, 54, 55, 56]. The introduction of the calibration step (by calculating the lower/upper bound ratios captured in Table 5) was to guide the decision of the final label prediction and avoid such an issue. The table demonstrates a clear separation of lower/upper bound ratios. Therefore, we further utilize such a mechanism by incorporating a calibrating step to further guide the classification process without having to train the algorithm with too many samples. The algorithmic steps for the calibration process are explained in Algorithm 4. Though the ranges provide an extra net for predicting the label, it is also possible that some document instances may fall outside the specified ranges of the datasets, which could result in not predicting a label correctly. Therefore, we introduced a proximity heuristic that favors the shortest distance to the ranges driven from the individual datasets (real or ChatGPT-generated) and assigns a label accordingly. Equations 5 and 6 demonstrate how the distance is calculated.
@@
Algorithm 5 illustrates the computational steps for multi-mode execution, demonstrating the complexity involved, including the proximity distance. To use the algorithm in detecting fake science, it must be trained using two different types of data: (1) a real publication dataset and (2) ChatGPT-generated articles. The algorithm also expects the ratio means of each data source, which are computed using the calibration algorithm.
@@
3:Input: a dataset $D$ of mixed fake and real documents
@@ 5 Discussion
In a world where generative AI has become widespread, various studies aimed to investigate the potential issues of using ChatGPT to generate fake science. The literature review showed a desperate need to advance the algorithmic approaches to discern real publications from fake ones, especially, when they are mixed. Our study aimed to address such issues incrementally. Specifically, we first tested the intuition of whether the content generated by ChatGPT may exhibit unique characteristics that distinguish it from real science. We explored this task using prompt engineering, where we created engineered datasets on the subjects of Alzheimer’s, cancer, and depression diseases. In this work, we contributed a prompt-engineering algorithm on how to generate simulated content to evaluate this premise. Working with plain text (using publication abstracts or generated from ChatGPT), using the TF-IDF algorithm is a common approach to generate bigrams that can be used to construct more complex models.
@@
Our initial observation of networks generated from ChatGPT content is that they are highly connected and contain fewer nodes compared to networks constructed from real publication text. Additionally, when we calculated the ratios of the number of bigrams against the total number of words of documents on k-Folds, we found that the ratios of ChatGPT content are much higher than scientific abstracts. These two indications supported our intuition that ChatGPT documents exhibit distinguishable behavior than PubMed abstracts. One interpretation of this observation could be due to the inherent design of the ChatGPT engine. As observed, ChatGPT is optimized to generate highly convincing content by predicting the next correlated terms statistically using a Large Language Model. On the other hand, scientists prioritize accurate documentation of hypothese testing, scientific experiments, and careful explanation of observations. Describing science in terms of highly correlated words is not a goal of scientists. Clearly, the difference in goals may contribute to less connectivity in scientific publications.
@@
Further, we introduced the xFakeSci algorithm, a learning algorithm that predicts a label for a given article. In the Methods section, we showed that it is designed to operate in two modes: (1) Single-mode: where only one type of articles from the same source is used for training and a new set of documents from the same pool is used for predicting the label of an article; and (2) Multi-mode: where the algorithm was trained from two sources and a hybrid train model (of real and generated datasets) was constructed to make the predictions. The single-model is trivial; therefore, we focused our experiments on demonstrating the multi-mode. We performed several experiments to do the following: (1) to test and measure how the xFakeSci algorithm predicts labels of ChatGPT generated documents for a given disease when mixed with scientific abstracts, (2) to evaluate whether the algorithm performs consistently using various datasets of different diseases not only one disease, (3) to test whether the year of the publication plays a role in predicting ChatGPT generated documents when mixed with publications from various periods (2020-2024, 2015-2019, 2010-2014), and (4) to benchmark the algorithm against a baseline of some of the most common data mining algorithms. Our results for each experiment used the TP, TN, FP, FN metrics and F1 scores.
@@
When testing whether the year of publication plays a role in label prediction, we observed F1 scores of 91%, 92%, and 94% for cancer-related publications across different periods. This suggests a pattern of better detection of ChatGPT articles when mixed with older publications. However, identifying newer publications proved more challenging. For the Alzheimer’s disease, while no improvement was observed, degradation was also absent. As mentioned earlier, the Alzheimer’s datasets were the smallest among all datasets, limiting the calibration process due to fewer k-Folds compared to other diseases. In the case of depression, the algorithm exhibited consistent performance with an F1 score of 80% across all periods. It’s plausible that mental health data acquisition posed limitations, potentially constraining resources from this specific area. Testing this hypothesis involves measuring document similarity between PubMed and ChatGPT sources using lexical and semantic analysis.
@@
Upon benchmarking xFakeSci against classical data mining algorithms, we observed an interesting pattern: xFakeSci correctly predicted all the scientific publications in all the experiments we performed. However, other algorithms misclassified publications as ChatGPT and vice versa (true positives, false positives, false negatives, and true negatives). xFakeSci, however, needed improvement in predicting true negatives (ChatGPT documents), as many ChatGPT documents were labeled as true positives (real publications). In all the experiments, the F1 scores of xFakeSci ranged between 80% and 94%. In contrast, the other data mining algorithms showed much lower performance, with F1 scores ranging between 32% and 52%. We attribute the high performance of xFakeSci to the calibration process, which was guided by ratios and proximity distances. Although the training model remained lightweight, both heuristics provided more guidance for predicting fake articles. This novel calibration method benefits from an abundance of data, without suffering from overfitting issues like other common classification algorithms. Clearly, the xFakeSci algorithm does not suffer from such a deficiency in identifying real articles when mixed with ChatGPT-generated content
@@ 6 Conclusions and Future Directions
When we asked a high school student about their knowledge of ChatGPT, they responded, “Do you mean that tool that does my homework for you?” Indeed, ChatGPT is an incredibly sophisticated tool with a wide range of impressive capabilities. Since the rise of ChatGPT, many new research topics have opened a new generative door, and many long-standing questions are now being investigated. However, the most significant concern associated with ChatGPT and other generative AI tools is that they could pose a threat to the future of science. If younger generations utilize ChatGPT to plagiarize, it could undermine the integrity of research and learning, potentially having a negative impact on the development of future pioneers.
@@
Looking ahead, there are several avenues for future research based on our current work: (1) conducting a preprocessing step (e.g., clustering) to group more closely related publications together (e.g., breast cancer, prostate cancer, and others), or separate diseases from co-morbidities. The use of knowledge graphs may be a powerful tool to use in continuing to investigate this research direction; (2) further experimentation in training and calibrating the xFakeSci algorithm by utilizing heuristics learned from preprocessing steps and the discoveries of clusters; and (3) testing the algorithm on more than two data sources (clinical reports, publications, and ChatGPT-generated documents).
@@ Acknowledgments
This research is supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement Sano No 857533 which is carried out within the International Research Agendas programme of the Foundation for Polish Science, co-financed by the European Union under the European Regional Development Fund, and created as part of the Ministry of Science and Higher Education’s initiative to support the activities of Excellence Centers established in Poland under the Horizon 2020 program based on the agreement No “MEiN/2023/DIR/3796”, and the National Natural Science Foundation of China (NSFC) under grant 62120106008. The authors also acknowledge Laila Hamed for her valuable perspective on ChatGPT.
@@
A.H. conceived the idea(s), A.A. and X.W. designed the experiment(s), A.A. and X.W. analyzed the results. Both authors written and reviewed the manuscript.
@@
Table 7: Multi-Mode Experiments: F1 Classification Scores From Most Recent Publications
@@
Table 8: Multi-Mode Experiments: F1 Classification Scores By Older Periods
@@
Figure 4: Multi-Mode Experiments: F1 Classification Scores for (Cancer, Alzheimer’s, and Depression) for publications gathered in period (2020-2024.)
Figure 5: Showing (TP, TN, FP, FN) Mertics and F1 Scores for xFakeSci vs Classical Data Mining algorithms Using Cancer Dataset From 2020-2024
Figure 6: Accuracy of xFakeSci vs Classical Data Mining Algorithms Showing Publication Year as a Prediction Factor.